the sum of 2 numbers is 12, and their product is 35. Find the numbers.
Answer by jim_thompson5910(28536)

"the sum of 2 numbers is 12" ---> x+y = 12 ---> y = 12-x
----------------------------

"their product is 35" ----> xy = 35

xy = 35

x(12-x) = 35

-x^2 + 12x = 35

-x^2 + 12x - 35 = 0

x^2 - 12x + 35 = 0

(x - 7)(x - 5) = 0

x-7 = 0 or x-5 = 0

x = 7 or x = 5

So the two numbers are 5 and 7
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------